# Talk2PC: Enhancing 3D Visual Grounding through LiDAR and Radar Point Clouds Fusion for Autonomous Driving

### v1

@@ III-D Dynamic Gated Graph Fusion
Secondly, to capture the regions of interest within the point cloud encoded by the textual prompt, as well as the surrounding contextual information, we leverage the power of graph structures, which excel at handling data with complex spatial or structural relationships. Given that point cloud data lacks a fixed grid structure, we model the features as a set of nodes within a graph. This approach enhances the ability to model contextual information, thereby improving the understanding of both the spatial distribution and the semantic information of the object. As Fig. 5 shows, compared with the static graph constructed by Sparse Vision Graph Attention (SVGA) [53], which cannot be dynamically adjusted according to the characteristics of different features, we construct the dynamic graph which leverages the axial structure of SVGA to maintain its efficiency while dynamically constructing a more representative graph, changing across the features. To achieve this, we estimate the mean ($\mu$) and standard deviation ($\sigma$) of the Euclidean distances between nodes by utilizing a node subset. This subset is obtained by partitioning the pseudo-image feature map of the point cloud into quadrants and comparing the diagonal pairs of these quadrants. This approach enables the computation of the estimated $\mu$ and $\sigma$ between the feature maps of the original pseudo-image and its diagonally flipped counterpart. The purpose of this estimation is to reduce the number of pairwise comparisons required to obtain $\mu$ and $\sigma$. We avoid directly computing the exact values of $\mu$ and $\sigma$ because doing so would necessitate calculating the Euclidean distance between every individual node and all other nodes in the image, which is computationally expensive. Subsequently, we adopt the row and column-wise connections as in SVGA to further reduce computational complexity. As demonstrated by MobileViG [53], every node should not be considered. If the Euclidean distance between two nodes is less than the difference between the estimated $\mu$ and $\sigma$, we connect the two nodes. DGGF also supports a variable number of connections across different point cloud feature maps, which contrasts with the standard approach in graph neural networks where a fixed number of K Nearest Neighbors (KNN) is used for all images. This is because, within different feature maps, the Euclidean distance between nodes is often smaller than the estimated difference between $\mu$ and $\sigma$. The intuition behind using $\mu$ and $\sigma$ is that nodes within the range of $\mu$$\pm$ $\sigma$ are close to each other and should share information. These values are then used to establish connections between the generated nodes, as shown in Figure 2b. In Fig. 5, we can observe that GGF connects the Region of Interest (RoI) of the object to non-object regions in the image, whereas our proposed DGGF only connects the object region to other object regions in the point cloud feature map.
@@
1:$\textit{X}_{norm}$leftarrow $\mathtt{norm}$(X, Xquadrants)   $\triangleright$ matrix norm of tensors
2:$\mu$leftarrow $\mathtt{mean}$($\textit{X}_{norm}$)
3:$\sigma$leftarrow $\mathtt{std}$($\textit{X}_{norm}$)
4:while $mK$ $<$ H do
5:  Xrolled $\leftarrow$ $\mathtt{roll}_{down}$(X, $mK$)
6:  dist $\leftarrow$ $\mathtt{norm}$(X, Xrolled)   $\triangleright$ get distance value
7:  if dist $<$ ($\mu$- $\sigma$) then   $\triangleright$ generate mask
8:    mask $\leftarrow$ 1
9:  else
10:    mask $\leftarrow$ 0
11:  end if
12:  Xmasked $\leftarrow$ mask * (Xrolled -- X)   $\triangleright$ get features
13:  Xfinal $\leftarrow$ $\mathtt{max}$(Xmasked, Xfinal)   $\triangleright$ keep max
14:  $m$leftarrow $m+1$
15:end while
16:$m$leftarrow 0
17:while $mK$ $<$ W do
18:  Xrolled $\leftarrow$ $\mathtt{roll}_{right}$(X, $mK$)
19:  dist $\leftarrow$ $\mathtt{norm}$(X, Xrolled)
20:  if dist $<$ ($\mu$- $\sigma$) then
21:    mask $\leftarrow$ 1
22:  else
23:    mask $\leftarrow$ 0
24:  end if
25:  Xmasked $\leftarrow$ mask * (Xrolled -- X)
26:  Xfinal $\leftarrow$ $\mathtt{max}$(Xmasked, Xfinal)
27:  $m$leftarrow $m+1$
28:end while
29:return $\mathtt{Conv2d}$($\mathtt{Concat}$(X, Xfinal))
@@
Through our proposed approach, DGGF constructs a more representative graph structure compared to GGF [37], as it does not connect dissimilar patches (i.e., nodes). Additionally, compared to KNN, DGGF significantly reduces the computational overhead by minimizing adjacency computations during graph construction (whereas KNN must determine the nearest neighbors for each image patch). Furthermore, DGGF does not require the reshaping necessary for performing graph convolution in KNN-based methods [54]. Therefore, DGGF combines the representational flexibility of KNN with the computational efficiency of GGF.
@@ III-E C3D-RECHead
Conventional 3D object detection models based on point clouds, such as CenterPoint [55], typically regress the bounding box relative to the object’s center. However, in 3D visual grounding based on point cloud sensors, we identify two key limitations of the center-based regression: (1) Higher Point Cloud Density at the Nearest Edge: The nearest object to the autonomous vehicle has a denser point cloud, especially on the side facing the vehicle, which provides a more reliable anchor point for regression, improving localization accuracy. (2) Alignment with Textual Distance Prompts: Many textual prompts in 3D visual grounding include object distances (e.g., ”the car 5 meters ahead”). These distances typically refer to the edge of the object closest to the vehicle, rather than the center. Thus, regressing from this nearest edge point aligns better with semantic and perceptual grounding.
Instead of regressing from the object’s center, we propose C3D-RECHead. Exactly, it performs bounding box regression from the nearest edge point to the vehicle. This adaptation preserves the computational efficiency of CenterPoint while improving localization accuracy in 3D visual grounding tasks. Assume a 3D bounding box $B$ is represented as:
@@ III-F Loss Function
Our proposed C3D-RECHead first generates heatmaps to predict the nearest corner location of each detected object, serving as an anchor point for subsequent regression. This approach leverages the increased density of LiDAR and radar point clouds around the closest object, improving localization accuracy. Following the heatmap-based localization, we refine key object properties under the CenterPoint-inspired supervision paradigm. Specifically, our model predicts:
  1. 1.
    Sub-voxel location refinement to enhance localization precision beyond the voxel grid resolution.
  2. 2.
    Height above ground to capture elevation information crucial for differentiating objects in 3D space.
  3. 3.
    3D bounding box dimensions include width, length, and height to characterize object shape and size.
  4. 4.
    Orientation angle to accurately model object rotation in the 3D scene.
@@
where ${L}_{\rm hm}$ is the confidence loss supervising the heatmap quality of the center-based detection head using a focal loss [56]; $\boldsymbol{\Lambda}=\left\{x,y,z,l,h,w,\theta\right\}$ indicates the smooth-$\ell_{1}$ loss supervising the regression of the box center (for slight modification based on heatmap peak guidance), dimensions, and orientation; and $\beta$ is the weight to balance the two components of the loss, which is set to 0.25 by default.
@@ IV Experiments and Performance Analysis
Table I: Overall performances on Talk2Radar dataset [37] (Best Performance. Car, Pedestrian and Cyclist provide specialized mAPs. Radar5 denotes the accumulation of five frame radar data.)
Models Sensors Text Encoder Fusion Entire Annotated Area (EAA) Driving Corridor Area (DCA)
@@
PointPillars Radar5 ALBERT [57] HDP 18.92 9.79 12.47 13.73 12.91 39.20 10.25 14.93 21.46 20.19
CenterFormer Radar5 ALBERT HDP 17.26 6.79 9.27 11.11 10.79 19.56 9.13 12.03 13.57 13.02
CenterPoint Radar5 ALBERT HDP 18.98 5.30 14.96 13.08 12.20 40.53 8.57 15.66 21.59 20.25
PointPillars Radar5 ALBERT MHCA 5.18 5.76 6.63 5.86 3.58 13.34 4.36 8.79 8.83 7.66
CenterFormer Radar5 ALBERT MHCA 4.53 3.48 4.00 4.00 2.03 8.77 3.52 6.69 6.33 5.92
CenterPoint Radar5 ALBERT MHCA 5.21 4.57 5.13 4.97 3.12 12.70 4.07 7.70 8.16 7.51
MSSG [36] Radar5 GRU [58] - 12.53 5.08 8.47 8.69 7.03 18.93 7.88 9.40 12.07 11.67
AFMNet [59] Radar5 GRU - 11.98 6.87 9.16 9.34 7.72 18.62 8.21 10.06 12.30 11.79
MSSG [36] Radar5 ALBERT - 16.03 5.86 10.57 10.82 8.96 25.79 8.69 12.55 15.68 14.12
AFMNet [59] Radar5 ALBERT - 16.31 6.80 10.35 11.15 9.46 26.82 8.71 12.45 15.99 14.18
EDA [60] Radar5 RoBERTa [50] - 13.23 6.60 8.63 9.49 8.93 23.55 8.80 11.95 14.77 13.07
T-RadarNet [37] Radar5 ALBERT GGF 24.68 9.71 15.74 16.71 14.88 42.58 10.13 17.82 23.51 22.37
TPCNet (ours) Radar5 PointCLIP [48] DGGF 25.92 9.60 16.37 17.30 15.29 45.97 12.73 19.73 26.14 24.73
CenterPoint LiDAR ALBERT HDP 28.16 6.21 17.46 17.28 16.03 43.43 6.87 27.18 25.83 24.93
CenterPoint LiDAR ALBERT MHCA 6.56 5.04 5.33 5.64 4.86 13.60 4.52 7.32 8.48 7.89
MSSG [36] LiDAR GRU - 15.38 7.52 11.67 11.52 9.76 23.27 8.68 13.51 15.15 14.75
AFMNet [59] LiDAR GRU - 16.13 7.68 12.51 12.11 9.92 24.50 9.07 13.87 15.81 15.11
MSSG [36] LiDAR ALBERT - 18.19 7.66 11.63 12.49 10.91 29.61 10.98 14.66 18.42 16.23
AFMNet [59] LiDAR ALBERT - 19.50 7.92 13.56 13.66 12.18 31.68 9.23 18.90 19.94 17.59
EDA [60] LiDAR RoBERTa - 16.10 6.91 12.88 11.96 10.10 25.10 9.28 15.73 16.70 14.91
T-RadarNet [37] LiDAR ALBERT GGF 24.91 12.74 18.67 18.77 17.20 48.98 14.69 27.24 30.30 29.89
TPCNet (ours) LiDAR PointCLIP DGGF 28.93 12.83 20.72 20.83 18.94 49.78 15.25 29.11 31.38 30.14
TPCNet (ours) LiDAR+Radar5 PointCLIP DGGF 32.13 15.99 23.74 23.95 22.01 52.72 17.30 31.66 33.89 31.73
@@ IV-A Settings of Models and Implementations
Model Settings: Besides the PointCLIP [48] that we leverge in our proposed TPCNet, for other models with pre-trained transformer (e.g., ALBERT [57]) for text encoding, we set the token length as 30, uniformly. The number of pillars in the pillar backbone of TPCNet is set as 10 and 32 for radar and LiDAR, respectively. We set the sizes of both LiDAR and Radar Context Agent Feature as $12\times 12$ in default.
@@
Dataset Settings: We conduct comprehensive training and evaluation of models on the Talk2Radar dataset [37], which encompasses three distinct categories of traffic participants: Car, Cyclist, and Pedestrian, utilizing both 4D radar and LiDAR modalities. Furthermore, to rigorously assess the generalization capabilities of T-RadarNet, we extend our evaluation to the Talk2Car dataset [33]. This dataset, derived from the nuScenes benchmark [67], provides synchronized LiDAR and radar data accompanied by textual prompts referencing object objects, thereby enabling a robust validation of our approach across diverse multi-modal scenarios.
@@
Models Sensors BEV AP 3D AP
[…]
Baseline [33] L 30.6 24.4 27.9 19.1
MSSG [36] L 27.8 26.1 31.9 20.3
EDA [60] L 37.0 29.8 37.2 20.4
AFMNet [59] L 45.3 33.1 41.9 20.7
T-RadarNet [37] L 52.8 39.9 47.2 30.5
TPCNet (ours) L+R 56.7 44.1 52.3 33.6
Training and Evaluation Settings: For the Talk2Radar dataset, all models are trained using a distributed setup across four RTX A4000 GPUs, with a batch size of 4 per GPU for a total of 80 epochs. The optimization process employs AdamW with an initial learning rate of $1\times 10^{-3}$, modulated by a cosine annealing learning rate scheduler, and a weight decay of $5\times 10^{-4}$. To evaluate 3D visual grounding performance, we utilize two key metrics: Average Precision (AP) and Average Orientation Similarity (AOS), which are computed for both the entire annotated area and the driving corridor. Specifically, we report the mean Average Precision (mAP) and mean Average Orientation Similarity (mAOS) for 3D bounding box localization and orientation estimation, respectively.
@@
IV-B Quantitative Results
Overall comparison on Talk2Radar and Talk2Car datasets. As shown in Table I, our proposed TPCNet, which integrates LiDAR and radar, achieves state-of-the-art performance on the Talk2Radar dataset. Compared to unimodal models (LiDAR-only and radar-only), TPCNet demonstrates a significant improvement in both mAP and mAOS metrics. Furthermore, even in its unimodal configuration, TPCNet outperforms the models with similar sensor inputs, highlighting the superiority of our model architecture. In terms of per-class accuracy, the detection performance for Car is slightly higher than that for Cyclist, while Pedestrian exhibits the lowest accuracy. Additionally, on the Talk2Car dataset, TPCNet also surpasses the best-performing LiDAR-only model, demonstrating strong generalization across sensor data distributions.
@@ IV-B Quantitative Results
Methods Talk2Radar Talk2Car
@@
Bi-MHCA [52] 29.14 25.87 49.7 33.2
Bi-MHLCA [65] 24.53 23.92 48.1 30.9
InterRAL [66] 26.45 24.23 50.8 31.3
L2R Fusion [27] 25.26 22.94 48.5 30.1
AGN [24] 27.84 25.03 50.1 31.0
BACA (ours) 28.92 26.07 52.3 33.6
Table V: Comparison of efficiency on fusion methods of LiDAR and radar features with different sizes. $N_{h}$ denotes the number of heads in cross attention, which is only for the attention-based fusion method. FPS is the frame per second of TPCNet equipped with various fusion methods on single RTX A4000 GPU.
@@
Stage 3 of pseudo image: $D$=64, $H$=80, $W$=80, $N_{h}$=4
Stage 4 of pseudo image: $D$=128, $H$=40, $W$=40, $N_{h}$=8
Stage 5 of pseudo image: $D$=256, $H$=20, $W$=20, $N_{h}$=8
@@
=18$\times$18 2.56K 27.04M
Performances of TPNet based on two input sensors for different types of prompts. Table III presents the performance of TPCNet with different sensor inputs when various types of prompts are processed. Overall, we observe that for text prompts containing motion or velocity information, radar generally outperforms LiDAR. In contrast, for prompts involving object depth, the LiDAR-only model performs better than the radar-only model. However, the fusion of radar and LiDAR significantly outperforms both unimodal approaches. Additionally, aggregating five frames of radar point clouds yields better performance than using three or one frame. This demonstrates the strong complementarity between radar and LiDAR.
Table VI: Comparison of performances between C3D-RECHead and CenterPoint-based prediction head.
@@
Performances of various fusion methods of LiDAR and radar under the settings of TPCNet. As TABLE IV shows, under the TPCNet architecture, our proposed BACA outperforms other LiDAR-radar fusion methods. Overall, it can be observed that the global fusion methods based on cross-attention achieve superior performance when compared with convolution-based fusion approaches. Specifically, the Bi-MHCA-based fusion method achieves the highest mAP, followed closely by our proposed BACA, which performs slightly worse than Bi-MHCA but surpasses Bi-MHLCA. Moreover, on the Talk2Car dataset, our BACA achieves the best performance.
Table VII: Statistics of mAP for predicted objects by TPCNet with the respect with depth upon 5-frame radar, LiDAR and the combination of radar and LiDAR.
Sensors Objects 0-10 (m) 10-20 20-30 30-40 40-50 50+
Radar5 Car 43.62 45.19 25.32 20.13 11.28 2.20
Pedestrian 16.79 8.11 3.98 3.09 4.89 0.0
Cyclist 31.55 14.23 16.17 5.74 3.23 1.22
LiDAR Car 46.78 49.50 31.72 26.45 14.19 8.12
Pedestrian 19.55 9.79 7.88 6.57 5.48 0.0
Cyclist 36.92 17.63 20.11 7.53 6.11 3.74
LiDAR Car 51.67 54.69 36.55 29.87 18.56 10.59
+ Pedestrian 21.41 11.16 7.92 7.10 5.79 0.0
Radar5 Cyclist 38.55 21.27 22.79 10.57 6.20 4.12
The comparison of efficiency between various fusion methods for LiDAR and radar. As shown in Table IX, we have compared the complexity and inference speeds of non-attention-based fusion methods (convolution-based AGN) and attention-based fusion methods for processing multi-scale feature maps from the two modalities. We observe that our proposed BACA has significantly fewer parameters than other fusion methods, with a 2 to 3 orders of magnitude reduction in the parameter count when compared to the other three methods. Moreover, the FLOPs of BACA are substantially lower than those of Bi-MHLCA, Bi-MHCA, and AGN. At the same time, BACA achieves a much higher FPS than other attention-based fusion methods and is very close to the convolution-based AGN in terms of inference speed. Furthermore, when setting the agent feature map size to 8×8, 16×16, and 18×18, the increase in FLOPs remains relatively small, demonstrating the strong scalability of BACA.
Table VIII: Ablation experiments of TPCNet with various settings on Talk2Radar dataset.

### v2

@@ III-D Dynamic Gated Graph Fusion
Secondly, to effectively capture regions of interest (RoIs) within the point cloud based on the textual prompt, while preserving surrounding contextual information, we employ graph-based modeling, which is well-suited for handling complex spatial and structural relationships. Since point cloud data lacks a fixed grid structure, we represent features as a set of nodes within a graph, enhancing the capability to model both spatial distributions and semantic attributes of objects.
As illustrated in Fig. 6, compared to the static graph constructed in Sparse Vision Graph Attention (SVGA) [29], which remains unchanged across different features, we propose a dynamic graph that retains SVGA’s axial structure efficiency while dynamically adjusting graph connectivity based on feature characteristics. To achieve this, rather than computing all pairwise distances exhaustively, we estimate the mean ($\mu$) and standard deviation ($\sigma$) of Euclidean distances between nodes by analyzing a subset of nodes, which is derived by partitioning the image into quadrants and evaluating the pairwise correspondence between diagonally opposite regions. Specifically, we partition the pseudo-image feature map of the point cloud into quadrants and compare diagonal pairs within these quadrants. This process provides an efficient approximation of $\mu$ and $\sigma$, significantly reducing the computational cost associated with exhaustive distance calculations.
Direct computation of exact $\mu$ and $\sigma$ values would require calculating Euclidean distances between every node pair, causing excessive computational complexity. Instead, we approximate these values using the aforementioned subset analysis, which balances efficiency and accuracy. Subsequently, we adopt row and column-wise connections, following SVGA, to further reduce computational overhead. As demonstrated by MobileViG [29], not all nodes need to be connected; instead, if the Euclidean distance between nodes is less than the estimated difference between $\mu$ and $\sigma$, a connection is established.
Unlike conventional GNN approaches, where a fixed number of K-nearest neighbors (KNN) is used for all images, our DGGF approach supports a variable number of connections across different point cloud feature maps. This flexibility allows adaptive connectivity, ensuring that node relationships better reflect the underlying feature distributions. The rationale for leveraging $\mu$ and $\sigma$ is that nodes within the range of $\mu\pm\sigma$ are likely to be spatially correlated and should exchange information. These values are subsequently used to establish graph connections, as illustrated in Fig. 6 (b).
As shown in Fig. 6, the GGF tends to connect RoIs with non-object regions, potentially introducing irrelevant information. In contrast, our proposed DGGF selectively connects object regions within the point cloud feature map, ensuring a more precise and contextually relevant feature representation.
@@
[…]
3:while mKmK << H do
4:  Xrolled $\leftarrow$ $\mathtt{roll}_{down}$(X, mKmK),  dist $\leftarrow$ $\mathtt{norm}$(X, Xrolled)   $\triangleright$ get distance value
5:  if dist << ($\mu$ - $\sigma$) thenmask $\leftarrow$ 1  $\triangleright$ generate mask
6:  elsemask $\leftarrow$ 0
7:  end if
8:  Xmasked $\leftarrow$ mask * (Xrolled - X)  $\triangleright$ get features,  Xfinal $\leftarrow$ $\mathtt{max}$(Xmasked, Xfinal)  $\triangleright$ keep max
9:  $m$ $\leftarrow$ $m+1$
10:end while
11:$m$ $\leftarrow$ 0
12:while mKmK << W do
13:  Xrolled $\leftarrow$ $\mathtt{roll}_{right}$(X, mKmK),  dist $\leftarrow$ $\mathtt{norm}$(X, Xrolled)
14:  if dist << ($\mu$ - $\sigma$) thenmask $\leftarrow$ 1
15:  elsemask $\leftarrow$ 0
16:  end if
17:  Xmasked $\leftarrow$ mask * (Xrolled - X),  Xfinal $\leftarrow$ $\mathtt{max}$(Xmasked, Xfinal)
18:  $m$ $\leftarrow$ $m+1$
19:end while
20:return $\mathtt{Conv2d}$($\mathtt{Concat}$(X, Xfinal))
@@
Through our proposed approach, DGGF constructs a more representative graph structure compared to GGF [14], as it does not connect dissimilar patches (i.e., nodes). Additionally, compared to KNN, DGGF significantly reduces the computational overhead by minimizing adjacency computations during graph construction (whereas KNN must determine the nearest neighbors for each image patch). Furthermore, DGGF does not require the reshaping necessary for performing graph convolution in KNN-based methods. Therefore, DGGF combines the representational flexibility of KNN with the computational efficiency of GGF.
@@ III-E C3D-RECHead
Conventional 3D object detection models based on point clouds, such as CenterPoint [30], typically regress the bounding box relative to the object’s center. However, in 3D visual grounding based on point cloud sensors, we identify two key limitations of the center-based regression: (1) Higher Point Cloud Density at the Nearest Edge: The nearest object to the autonomous vehicle has a denser point cloud, especially on the side facing the vehicle, which provides a more reliable anchor point for regression, improving localization accuracy. (2) Alignment with Textual Distance Prompts: Many textual prompts in 3D visual grounding include object distances (e.g., “the car 5 meters ahead”). These distances typically refer to the edge of the object closest to the vehicle, rather than the center. Thus, regressing from this nearest edge point aligns better with semantic and perceptual grounding.
Instead of regressing from the object’s center, we propose C3D-RECHead (Fig. 7). Exactly, it performs bounding box regression from the nearest edge point to the vehicle. This adaptation preserves the computational efficiency of CenterPoint while improving localization accuracy in 3D visual grounding tasks. Assume a 3D bounding box $B$ is represented as:
@@ III-F Loss Function
The proposed C3D-RECHead first generates heatmaps to estimate the nearest corner of each detected object, which serves as an anchor point for subsequent regression. This design exploits the higher point density from LiDAR and radar near the closest object surface, thereby enhancing localization accuracy. Conditioned on the heatmap localization, the head refines object attributes following a CenterPoint-inspired supervision scheme. Specifically, it predicts: (1) sub-voxel offsets for localization precision beyond the voxel resolution; (2) height above ground to capture vertical positioning critical for 3D discrimination; (3) 3D bounding box dimensions $(w,l,h)$ to describe object geometry; and (4) orientation to model rotational pose within the 3D scene.
@@
where ${L}_{\rm hm}$ is the confidence loss supervising the heatmap quality of the center-based detection head using a focal loss; $\boldsymbol{\Lambda}=\left\{x,y,z,l,h,w,\theta\right\}$ indicates the smooth-$\ell_{1}$ loss supervising the regression of the box center (for slight modification based on heatmap peak guidance), dimensions, and orientation; and $\beta$ is the weight to balance the two components of the loss, which is set to 0.25 by default.
@@ IV Experiments and Performance Analysis
Table I: Overall performances on Talk2Radar dataset [14] (Best Performance. Car, Pedestrian and Cyclist provide specialized mAPs. R5 denotes the accumulation of five frame radar data; C denotes camera while L denotes LiDAR.)
Models Venues Sensors Text Encoder Fusion Entire Annotated Area (EAA) Driving Corridor Area (DCA)
@@
[…]
TPCNet (ours) 2025 R5 PointCLIP [25] DGGF 25.92 9.60 16.37 17.30 15.29 45.97 12.73 19.73 26.14 24.73
[…]
TPCNet (ours) 2025 L PointCLIP DGGF 28.93 12.83 20.72 20.83 18.94 49.78 15.25 29.11 31.38 30.14
[…]
TPCNet (ours) 2025 L + R5 PointCLIP DGGF 32.13 15.99 23.74 23.95 22.01 52.72 17.30 31.66 33.89 31.73
@@ IV-A Settings of Models and Implementations
Model Settings: Besides the PointCLIP [25] that we leverge in our proposed TPCNet, for other models with pre-trained transformer (e.g., ALBERT [31]) for text encoding, we set the token length as 30, uniformly. The number of pillars in the pillar backbone of TPCNet is set as 10 and 32 for radar and LiDAR, respectively. We set the sizes of both LiDAR and Radar Context Agent Feature as $12\times 12$ in default. For the $K$ values of graph construction in DGGF, we set 8, 4, 2 for three stage feature maps, respectively.
@@
Dataset Settings: We conduct comprehensive training and evaluation of models on the Talk2Radar dataset [14], which encompasses three distinct categories of traffic participants: Car, Cyclist, and Pedestrian, utilizing both 4D radar and LiDAR modalities. Furthermore, to rigorously assess the generalization capabilities of T-RadarNet, we extend our evaluation to the Talk2Car dataset [11], which provides synchronized LiDAR and radar data accompanied by textual prompts referencing object objects, thereby enabling a robust validation of our approach across diverse multi-modal scenarios.
@@
Models Sensors BEV AP 3D AP Rain Night
$_{\text{A}}  \uparrow AP$$_{\text{B}}  \uparrow$ $_{\text{A}}  \uparrow AP$$_{\text{B}}  \uparrow$ $_{\text{A}}  \uparrow AP$$_{\text{A}}  \uparrow$
Baseline [11] L 30.6 24.4 27.9 19.1 - -
MSSG [13] L 27.8 26.1 31.9 20.3 - -
EDA [34] L 37.0 29.8 37.2 20.4 18.9 13.2
AFMNet [33] L 45.3 33.1 41.9 20.7 23.8 15.0
T-RadarNet [14] L 52.8 39.9 47.2 30.5 24.3 17.4
TPCNet (ours) L 53.3 42.7 46.8 31.6 24.9 19.0
TPCNet (ours) L+R 56.7 44.1 52.3 33.6 27.5 21.8
Training and Evaluation Settings: For the Talk2Radar dataset, all models are trained using a distributed setup across four RTX A4000 GPUs, with a batch size of 4 per GPU for a total of 80 epochs. The optimization process employs AdamW with an initial learning rate of $1\times 10^{-3}$, modulated by a cosine annealing learning rate scheduler, and a weight decay of $5\times 10^{-4}$. To evaluate 3D visual grounding performance, we utilize two key metrics: Average Precision (AP) and Average Orientation Similarity (AOS), which are computed for both the entire annotated area and the driving corridor. Specifically, we report the mean Average Precision (mAP) and mean Average Orientation Similarity (mAOS) for 3D bounding box localization and orientation estimation.
@@
IV-B Comparison with State-of-the-arts
Overall comparison with state-of-the-art approaches on Talk2Radar and Talk2Car datasets. As shown in Table I, our proposed TPCNet, which integrates LiDAR and radar, achieves state-of-the-art performance on the Talk2Radar dataset. Compared to unimodal models (LiDAR-only and radar-only), TPCNet demonstrates a significant improvement in both mAP and mAOS metrics. Furthermore, even in its unimodal configuration, TPCNet outperforms the models with similar sensor inputs, highlighting the superiority of our model architecture. In terms of per-class accuracy, the detection performance for Car is slightly higher than that for Cyclist, while Pedestrian exhibits the lowest accuracy. Additionally, on the Talk2Car dataset (Table II), TPCNet also surpasses the best-performing LiDAR-only model, demonstrating strong generalization across sensor data distributions.
@@ IV-B Comparison with State-of-the-arts
Table IV: Statistics of mAP for predicted objects by TPCNet with the respect with depth upon 5-frame radar, LiDAR and the combination of radar and LiDAR.
Sensors Objects 0-10 (m) 10-20 20-30 30-40 40-50 50+
R5 Car 43.62 45.19 25.32 20.13 11.28 2.20
Pedestrian 16.79 8.11 3.98 3.09 4.89 0.0
Cyclist 31.55 14.23 16.17 5.74 3.23 1.22
L Car 46.78 49.50 31.72 26.45 14.19 8.12
Pedestrian 19.55 9.79 7.88 6.57 5.48 0.0
Cyclist 36.92 17.63 20.11 7.53 6.11 3.74
L Car 51.67 54.69 36.55 29.87 18.56 10.59
+ Pedestrian 21.41 11.16 7.92 7.10 5.79 0.0
R5 Cyclist 38.55 21.27 22.79 10.57 6.20 4.12
Performances of TPNet based on two input sensors for different types of prompts. Table III presents the performance of TPCNet with different sensor inputs when various types of prompts are processed. Overall, for text prompts containing motion or velocity information, radar generally outperforms LiDAR. In contrast, for prompts involving object depth, the LiDAR-only model performs better than the radar-only model. However, the fusion of radar and LiDAR significantly outperforms both uni-modal approaches. Additionally, aggregating five frames of radar point clouds yields better performance than using three or one frame. This demonstrates the strong complementarity between radar and LiDAR.
@@
Methods Venues Talk2Radar Talk2Car
mAP  $\uparrow$ mAOS  $\uparrow$ $_{\text{A}}  \uparrow AP$$_{\text{B}}  \uparrow$
@@
[…]
BACA (ours) 2025 28.92 26.07 52.3 33.6
Performances of various fusion methods of LiDAR and radar under the settings of TPCNet. As TABLE V shows, under the TPCNet architecture, our BACA outperforms other LiDAR-radar fusion methods. Overall, it can be observed that the global fusion methods based on cross-attention achieve superior performance compared with convolution-based fusion approaches. Specifically, the Bi-MHCA-based fusion method achieves the highest mAP, followed closely by our proposed BACA, which performs slightly worse than Bi-MHCA but surpasses Bi-MHLCA. Moreover, on the Talk2Car dataset, our BACA achieves the best performance.
Table VI: Comparison of efficiency on fusion methods of LiDAR and radar features with different sizes. $N_{h}$ denotes the number of heads in cross attention, which is only for the attention-based fusion method. FPS is the frame per second of TPCNet equipped with various fusion methods on single RTX A4000 GPU.
@@
Stage 3 of pseudo image: DD=64, HH=80, WW=80, $N_{h}$=4
Stage 4 of pseudo image: DD=128, HH=40, WW=40, $N_{h}$=8
Stage 5 of pseudo image: DD=256, HH=20, WW=20, $N_{h}$=8
@@
$\times18$ 2.56K 27.04M
The comparison of efficiency between various fusion methods for LiDAR and radar. As shown in Table VI, we have compared the complexity and inference speeds of non-attention-based fusion methods (convolution-based AGN) and attention-based fusion methods for processing multi-scale feature maps from the two modalities. We observe that our proposed BACA has significantly fewer parameters than other fusion methods, with a 2 to 3 orders of magnitude reduction in the parameter count when compared to the other three methods. Moreover, the FLOPs of BACA are substantially lower than those of Bi-MHLCA, Bi-MHCA, and AGN. At the same time, BACA achieves a much higher FPS than other attention-based fusion methods and is very close to the convolution-based AGN in terms of inference speed. Furthermore, when setting the agent feature map size to 8×8, 16×16, and 18×18, the increase in FLOPs remains relatively small, demonstrating the strong scalability of BACA.
The comparison between CenterPoint-based prediction head and our proposed C3D-RECHead. As shown in Table VII, we comprehensively evaluate the accuracy of 3D object detection and orientation angle estimation using two different detection head paradigms on the Talk2Radar dataset. Our findings indicate that the proposed C3D-RECHead exhibits a significant performance advantage over CenterPoint, achieving notable improvements in both mAP and mAOS.
Table VII: Comparison of performances between C3D-RECHead and CenterPoint-based prediction head.
@@
[…]
Prompt Motion Depth Velocity
Methods Car Pedes Cyclist Car Pedes Cyclist Car Pedes Cyclist
[…]
BACA 38.9 14.1 21.2 49.7 25.7 42.7 40.8 28.6 32.0
Figure 8: Visualization of predicted results by proposed TPCNet in Talk2Radar dataset. The first row presents the prediction from the view of camera while the second row shows the results from Bird’s Eye View (BEV).
The comparison of the architecture of models. As Fig. 10 shows, we have compared T-RadarNet and TPCNet, with the models taking radar and LiDAR as single-modal inputs, respectively, to analyze the effectiveness of the model architecture. Our results show that, regardless of whether radar or LiDAR is used as the input, TPCNet consistently outperforms T-RadarNet in accuracy across all object categories and at various distance intervals from the ego vehicle, demonstrating the architectural superiority of TPCNet over T-RadarNet.
Statistics of prediction accuracy at different depths. We partition the predicted objects into six depth intervals, each spanning 10 meters relative to the ego vehicle. As shown in Table IV and Fig. 9, the fusion of two sensors significantly enhances object detection accuracy, particularly for distant objects. The fused model outperforms both LiDAR-only and radar-only modalities, demonstrating the necessity of multi-modal fusion.
Table IX: Ablation experiments of TPCNet with various settings on Talk2Radar dataset.